If you have $3000 and you want to deposit it into a savings account that offers .10% apr, how much would the amount be if you leave it there for six years? Assume that it's compounded monthly.
Answer #1

Answer)

Amount = P*(1+r)^n

P = 3000

r = 10/1200 {as there are 12 months in a year}

n = 6*12 = 72

After substitution

Amount = 5452.78284068